# Closing the Gap on the Sample Complexity of 1-Identification

### v1

@@ A.1 Simplification of History Result
A.1.1 Theorem 2 in Poiani et al. (2025)
Theorem 2 in Poiani et al. (2025) presents the Non-asymptotic analysis for algorithm S-TaS. For an instance $\nu$ with $r_{a}^{\nu}=\mu_{a}$, $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and , they define $\neg a=\{\nu^{\prime}:r_{a}^{\nu^{\prime}}<\mu_{0}\}$ for $a\in[m]$, and
@@
Then, Theorem 2 in Poiani et al. (2025) considers $\epsilon_{\nu}$ such that for all instance $\nu^{\prime}$ such that
@@
Thus, we can conclude $\frac{2\log(1/\delta)}{\Delta_{0,1}^{2}}+\frac{K^{2}}{\epsilon_{\nu}^{4}}$ is smaller than the proposed upper bound in Poiani et al. (2025). The remaining work is to show $\epsilon_{\nu}\leq\Delta_{0,1}$. Denote $a^{\text{sub}}=\arg\max_{a:\mu_{a}<\mu_{1}}\mu_{a}$. We have
@@
A.1.2 Theorem 3 in Kano et al. (2017)
Following Appendix A.2 in Li and Cheung (2025), we can also simplify the Theorem 3 in Kano et al. (2017). By applying the algorithm HDoC, for $\nu\in\mathcal{S}^{\text{pos}}$, we have
@@
A.1.3 Thereom 2 in Jourdan and Réda (2023)
Jourdan and Réda (2023) do not explicitly provide upper bound for $\mathbb{E}\tau$. Therorem 2 in Jourdan and Réda (2023) suggests that for a positive instance $\nu$.
@@ A.1.3 Thereom 2 in Jourdan and Réda (2023)
$W_{-1}$ is the negative branch of the Lambert W function, and $D^{\text{pos}}(\{\mu_{a}\}_{a=1}^{K})$ is further defined in the Appendix F in Jourdan and Réda (2023), as
@@
A.1.4 Theorem 8 in Katz-Samuels and Jamieson (2020)
As discussed in the Appendix A.1 in Li and Cheung (2025), proposed Algorithm FWER-FWPD only work on the positive instance with a different definition of $\mathbb{E}\tau$. But we can adapt their result. Apply Algorithm FWER-FWPD on a positive instance $\nu$, Theorem 8 in Katz-Samuels and Jamieson (2020) suggests that
@@ B.1 Proof of Preparation
The first step is to prove Lemma 4, showing that we suffice to only consider symmetric algorithm
@@
$\Pr_{\widetilde{\text{alg}},\sigma(\nu^{\prime})}(\forall t\in[T],A_{t}=a,\tau>t)=\frac{\sum_{\tilde{\sigma}\in S}\Pr_{\text{alg},\tilde{\sigma}(\nu^{\prime})}(\forall t\in[T],A_{t}=a,\tau>t)}{K!},$
which means $\tilde{\text{alg}}$ fits in the definition. Then, we have
@@
The next step, which is also the main trick, is to the prove Lemma 7.
@@
Given a pair $i,j\in[K],i\neq j$ and an arbitrary instance $\nu^{\prime}$, denote $\Delta=|r^{\nu^{\prime}}_{i}-r^{\nu^{\prime}}_{j}|$. Define
@@
An observation is given the realized reward and a symmetric algorithm, the set $\{t:N_{i}(t)+N_{j}(t)=n,A_{t+1}\in\{i,j\}\}=:S_{ij}(n)$ contains at most one element. If $|S_{ij}(n)|=1$, $\tau_{n}$ will be the unique element, and we have $\tau_{n}<\tau$ holds in this case. If $|S_{ij}(n)|=0$, we have $\tau_{n}=\infty$, meaning that $\tau_{n}\geq\tau$. Another observation is
@@
is the kl-divergence between two Bernoulli distribution with probability $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=i|\tau_{n}<\tau)$, $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=j|\tau_{n}<\tau)$, and not hard to see $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=i|\tau_{n}<\tau)+\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=j|\tau_{n}<\tau)=1$.
@@
Denote $\nu^{\prime\prime}$ as the instance which switch the position of arm $i$ and arm $j$ in $\nu^{\prime}$, i.e. $r_{i}^{\nu^{\prime}}=r_{j}^{\nu^{\prime\prime}},r_{j}^{\nu^{\prime}}=r_{i}^{\nu^{\prime\prime}}$ and $r_{a}^{\nu^{\prime}}=r_{a}^{\nu^{\prime\prime}},a\neq i,j$.
We first derive
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}\leq\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$ (15)
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}\leq\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$ (16)
where $Z_{s}|Z_{1},Z_{2},\cdots,Z_{s-1}\sim N(\frac{\Delta^{2}}{2},\Delta^{2})$. Since these two inequalities are symmetric, we suffice to prove the first one. Take $\mathcal{E}=\{\tau_{n}<\tau,A_{\tau_{n}+1}=i\}$, from the Lemma 18 in Kaufmann et al. (2016), we have
$\Pr_{\nu^{\prime\prime}}(\mathcal{E})=$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)$
$=$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})+$
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)|\neg\mathcal{E}\right]\Pr_{\nu^{\prime}}(\neg\mathcal{E})$
$=$ $\mathbb{E}_{\nu^{\prime}}\left[\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})$
$\geq$ $\exp\left(\mathbb{E}_{\nu^{\prime}}\left[-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}|\mathcal{E}\right]\right)\Pr_{\nu^{\prime}}(\mathcal{E}),$
@@
$\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}|\mathcal{E}\right]$
$=$ $\frac{\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)\right]}{\Pr_{\nu^{\prime}}(\mathcal{E})},$
and we can conclude $\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\geq\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)\right]$. From the above calculation, we know $\hat{KL}_{i,s},\hat{KL}_{j,s}$ both follow $N(\frac{(r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}})^{2}}{2},(r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}}))$, and we can rewrite the conclusion as
$\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right]$
@@
By the fact that the algorithm is symmetric, we can conclude $\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)=\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)$ and $\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)=\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)$. Sum up inequality (15) and (16), we get
@@
$\leq$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right).$
@@
$\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$
@@
Denote $D:=\max_{1\leq t\leq n}\sum_{s=1}^{t}\left(Z_{s}-\frac{\Delta^{2}}{2}\right)$, we have
$\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$
@@
The remaining work is to prove $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)D\leq\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)$. To simplify the notation, we take event $\mathcal{E}=\{\tau_{n}<\tau\}$ here and we can conduct the following calculation.
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})D\right]$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\mathds{1}(D\geq 0)D\right]$
$=$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathds{1}(\mathcal{E})\mathds{1}(D\geq 0)D>s\Big)ds$
$\leq$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathcal{E},D>s\Big)ds$
$\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}),\Pr_{\nu^{\prime}}\Big(D>s\Big)\right\}ds$
From Lemma 32, we can conclude $\Pr_{\nu^{\prime}}\big(D\geq s\big)\leq\exp(-\frac{s^{2}}{2n\Delta^{2}})$. Plug in the inequality, we get
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})D\right]\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}),\exp(-\frac{s^{2}}{2n\Delta^{2}})\right\}ds$
$=$ $\int_{0}^{\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}\Pr_{\nu^{\prime}}(\mathcal{E})ds+\int_{\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}^{+\infty}\exp(-\frac{s^{2}}{2n\Delta^{2}})ds$
$=$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\frac{\sqrt{2\pi}\sqrt{n\Delta^{2}}}{\sqrt{2\pi}}\int_{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}^{+\infty}\exp(-\frac{s^{2}}{2})ds$
$\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n\Delta^{2}}\exp(-\frac{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}{2})$
$=$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\mathcal{E}).$
@@
Proof.
From the Lemma 18 in Kaufmann et al. (2016), we have

### v2

@@ A.1 Simplification of History Result
A.1.1 Theorem 2 in Poiani et al. [2025]
Theorem 2 in Poiani et al. [2025] presents the Non-asymptotic analysis for algorithm S-TaS. For an instance $\nu$ with $r_{a}^{\nu}=\mu_{a}$, $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and , they define $\neg a=\{\nu^{\prime}:r_{a}^{\nu^{\prime}}<\mu_{0}\}$ for $a\in[m]$, and
@@
Then, Theorem 2 in Poiani et al. [2025] considers $\epsilon_{\nu}$ such that for all instance $\nu^{\prime}$ such that
@@
Thus, we can conclude $\frac{2\log(1/\delta)}{\Delta_{0,1}^{2}}+\frac{K^{2}}{\epsilon_{\nu}^{4}}$ is smaller than the proposed upper bound in Poiani et al. [2025]. The remaining work is to show $\epsilon_{\nu}\leq\Delta_{0,1}$. Denote $a^{\text{sub}}=\arg\max_{a:\mu_{a}<\mu_{1}}\mu_{a}$. We have
@@
A.1.2 Theorem 3 in Kano et al. [2017]
Following Appendix A.2 in Li and Cheung [2025], we can also simplify the Theorem 3 in Kano et al. [2017]. By applying the algorithm HDoC, for $\nu\in\mathcal{S}^{\text{pos}}$, we have
@@
A.1.3 Thereom 2 in Jourdan and Réda [2023]
Jourdan and Réda [2023] do not explicitly provide upper bound for $\mathbb{E}\tau$. Therorem 2 in Jourdan and Réda [2023] suggests that for a positive instance $\nu$.
@@ A.1.3 Thereom 2 in Jourdan and Réda [2023]
$W_{-1}$ is the negative branch of the Lambert W function, and $D^{\text{pos}}(\{\mu_{a}\}_{a=1}^{K})$ is further defined in the Appendix F in Jourdan and Réda [2023], as
@@
A.1.4 Theorem 8 in Katz-Samuels and Jamieson [2020]
As discussed in the Appendix A.1 in Li and Cheung [2025], proposed Algorithm FWER-FWPD only work on the positive instance with a different definition of $\mathbb{E}\tau$. But we can adapt their result. Apply Algorithm FWER-FWPD on a positive instance $\nu$, Theorem 8 in Katz-Samuels and Jamieson [2020] suggests that
@@ B.1 Proof of Preparation
The first step is to prove Lemma 4, showing that we suffice restrict attention to symmetric algorithms.
@@
$\Pr_{\widetilde{\text{alg}},\sigma(\nu^{\prime})}(\forall t\in[T],A_{t}=\sigma(a),\tau>t)=\frac{\sum_{\tilde{\sigma}\in S}\Pr_{\text{alg},\tilde{\sigma}(\nu^{\prime})}(\forall t\in[T],A_{t}=\tilde{\sigma}(a),\tau>t)}{K!},$
which means $\widetilde{\textnormal{alg}}$ fits in the definition. Then, we have
@@
The next step establishes a technical change-of-measure bound that controls a conditional KL divergence. This lemma is the main device used to argue that a symmetric algorithm cannot reliably distinguish two arms when the total number of samples from the pair is small.
@@
Fix a pair $i,j\in[K],i\neq j$ and an arbitrary instance $\nu^{\prime}$, denote $\Delta=|r^{\nu^{\prime}}_{i}-r^{\nu^{\prime}}_{j}|$. Define
@@
An observation is given a fixed sample path and a symmetric algorithm, the set $\{t:N_{i}(t)+N_{j}(t)=n,A_{t+1}\in\{i,j\}\}=:S_{ij}(n)$ contains at most one element. If $|S_{ij}(n)|=1$, $\tau_{n}$ will be the unique element, and $\tau_{n}<\tau$ holds; if $|S_{ij}(n)|=0$, then $\tau_{n}=\infty$ and hence $\tau_{n}\geq\tau$. Moreover,
@@
is the kl-divergence between two Bernoulli distribution with probability $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=i|\tau_{n}<\tau)$, $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=j|\tau_{n}<\tau)$. From the definition of $\tau_{n}$, not hard to see $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=i|\tau_{n}<\tau)+\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=j|\tau_{n}<\tau)=1$.
@@
Denote $\nu^{\prime\prime}$ as the instance which switch the position of arm $i$ and arm $j$ in $\nu^{\prime}$, i.e. $r_{a}^{\nu^{\prime\prime}}=\begin{cases}r_{j}^{\nu^{\prime}}&a=i\\ r_{i}^{\nu^{\prime}}&a=j\\ r_{a}^{\nu^{\prime}}&\text{else}\end{cases}$. We first derive
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}\leq\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\left(\sum_{s=1}^{N_{i}(\tau_{n})+N_{j}(\tau_{n})}Z_{s}\right)$ (13)
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}\leq\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\left(\sum_{s=1}^{N_{i}(\tau_{n})+N_{j}(\tau_{n})}Z_{s}\right)$ (14)
where $Z_{s}|Z_{1},Z_{2},\cdots,Z_{s-1}\sim N(\frac{\Delta^{2}}{2},\Delta^{2})$. Since these two inequalities are symmetric, we suffice to prove the first (13). Take $\mathcal{E}=\{\tau_{n}<\tau,A_{\tau_{n}+1}=i\}$. Notice that $\mathcal{E}$ is $\tau_{n}$-measurable, as both $\{\tau_{n}<\tau\},\{A_{\tau_{n}+1}=i\}$ are $\tau_{n}$-measurable. From the Lemma 18 in Kaufmann et al. [2016],
$\Pr_{\nu^{\prime\prime}}(\mathcal{E})=$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}\right)$
$=$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})+$
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}\right)|\neg\mathcal{E}\right]\Pr_{\nu^{\prime}}(\neg\mathcal{E})$
$=$ $\mathbb{E}_{\nu^{\prime}}\left[\exp\left(-\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})$
$\geq$ $\exp\left(\mathbb{E}_{\nu^{\prime}}\left[-\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}|\mathcal{E}\right]\right)\Pr_{\nu^{\prime}}(\mathcal{E}),$
@@
$\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}|\mathcal{E}\right]$
$=$ $\frac{\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}\right)\right]}{\Pr_{\nu^{\prime}}(\mathcal{E})},$
and we can conclude $\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\geq\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}\right)\right]$. From the above calculation, we know $\hat{KL}_{i,s},\hat{KL}_{j,s}$ both follow $N(\frac{(r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}})^{2}}{2},(r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}}))$, and we can rewrite the conclusion as
$\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{s=1}^{N_{i}(\tau_{n})+N_{j}(\tau_{n})}Z_{s}\right]$
@@
By symmetry of the algorithm, we can conclude $\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)=\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)$ and $\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)=\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)$. Sum up inequality (13) and (14), we get
@@
$\leq$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau_{n})+N_{j}(\tau_{n})}Z_{s}\right).$
@@
$\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau_{n})+N_{j}(\tau_{n})}Z_{s}\right)$
@@
Let $D:=\max_{1\leq t\leq n}\sum_{s=1}^{t}\left(Z_{s}-\frac{\Delta^{2}}{2}\right)$. Since $\{\tau_{n}<\tau\}$ implies $N_{i}(\tau_{n})+N_{j}(\tau_{n})\leq n$, we have
$\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau_{n})+N_{j}(\tau_{n})}Z_{s}\right)$
@@
The remaining work is to prove $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)D\leq\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)$. To simplify the notation, we take event $\mathcal{E}_{n}=\{\tau_{n}<\tau\}$ here and use the layer-cake representation,
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E}_{n})D\right]$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E}_{n})\mathds{1}(D\geq 0)D\right]$
$=$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathds{1}(\mathcal{E}_{n})\mathds{1}(D\geq 0)D>s\Big)ds$
$\leq$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathcal{E}_{n},D>s\Big)ds$
$\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}_{n}),\Pr_{\nu^{\prime}}\Big(D>s\Big)\right\}ds$
From Lemma 32, we can conclude $\Pr_{\nu^{\prime}}\big(D\geq s\big)\leq\exp(-\frac{s^{2}}{2n\Delta^{2}})$. Plugging this tail bound yields
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E}_{n})D\right]\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}_{n}),\exp(-\frac{s^{2}}{2n\Delta^{2}})\right\}ds$
$=$ $\int_{0}^{\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}}\Pr_{\nu^{\prime}}(\mathcal{E}_{n})ds+\int_{\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}}^{+\infty}\exp(-\frac{s^{2}}{2n\Delta^{2}})ds$
$=$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}\Pr_{\nu^{\prime}}(\mathcal{E}_{n})+\frac{\sqrt{2\pi}\sqrt{n\Delta^{2}}}{\sqrt{2\pi}}\int_{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}}^{+\infty}\exp(-\frac{s^{2}}{2})ds$
$\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}\Pr_{\nu^{\prime}}(\mathcal{E}_{n})+\sqrt{2\pi n\Delta^{2}}\exp(-\frac{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}}{2})$
$=$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}\Pr_{\nu^{\prime}}(\mathcal{E}_{n})+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\mathcal{E}_{n}).$
@@
Proof of Lemma 9.
From the Lemma 18 in Kaufmann et al. [2016], we have